OpenStudy (kelhina):

China's population is 14.46 *10^8. Marshall Islands population is 602.5*10^2. How many times larger is china population than marshalls islands population? Enter your answer using scientific notation

OpenStudy (mhchen):

This is basic subtraction. To find the difference between 2 numbers, you subtract them.

OpenStudy (kelhina):

It is definitely not subtraction

OpenStudy (kelhina):

To find the difference you actually divide

OpenStudy (kelhina):

@jackthegreatest can you help me (:

OpenStudy (mhchen):

Well it says "How many ONES larger" so if whatever that means. IF it means how MUCH larger, it's subtraction. IF it means how many TIMES larger, it's division.

OpenStudy (kelhina):

I actually messed up writing the question it's how many times larger

OpenStudy (kelhina):

@mhchen

OpenStudy (kelhina):

Can someone help me (:

OpenStudy (studygurl14):

\(\huge \frac{14.46 \times 10^8}{602.5\times 10^2} = \frac{14.46}{602.5}\times\frac{10^8}{10^2}=?\) Does that help?

OpenStudy (kelhina):

i know how to set it up im just confused on how to actually solve it

OpenStudy (jackthegreatest):

ok so you're confused on dividing 10^8 by 10^2?

OpenStudy (kelhina):

well not exactly wouldnt u have to change the powers 10^8 and 10^2 because 14.46 and 602.5 arent in scientfic notation

OpenStudy (jackthegreatest):

they do not need to be

OpenStudy (kelhina):

so then it would be 1.446 and 6.025 and the powers would be 9 and 4

OpenStudy (kelhina):

oh really ?

OpenStudy (kelhina):

then what do i do subtract the powers or add them

OpenStudy (studygurl14):

dividing means you subtract the powers. So \(\large 10^{8-2}\)

OpenStudy (jackthegreatest):

subtract

OpenStudy (jackthegreatest):

^^

OpenStudy (kelhina):

then what with the decimals just divide them just like that even with the decimal in that spot?

OpenStudy (studygurl14):

yes. Tell me what you get from 14.46/602.5

OpenStudy (kelhina):

so then it would be 0.024 right?

OpenStudy (kelhina):

so the answer is 0.0024 10^6

OpenStudy (studygurl14):

yep. So you get \(\Large 0.024 \times 10^{6}\) But there is still more to be done

OpenStudy (kelhina):

didnt mean to add 2 zeros

OpenStudy (studygurl14):

Now you need to conver 0.024 to scientific notation

OpenStudy (kelhina):

2.4 ?

OpenStudy (jackthegreatest):

yep, how many powers of 10 were needed

OpenStudy (studygurl14):

yes. 2.4 times what htough?

OpenStudy (kelhina):

2

OpenStudy (jackthegreatest):

yep

OpenStudy (studygurl14):

Almost. -2

OpenStudy (jackthegreatest):

10^6-2 ...

OpenStudy (kelhina):

oh okay

OpenStudy (studygurl14):

So You get \(\large 2.4\times 10^{6-2}\)

OpenStudy (jackthegreatest):

^^

OpenStudy (kelhina):

thats the answr or do i change it again haha

OpenStudy (kelhina):

10^4

OpenStudy (jackthegreatest):

10^4

OpenStudy (jackthegreatest):

well 2.4 x 10^4

OpenStudy (jackthegreatest):

yes

OpenStudy (kelhina):

well yeah haha of course like that

OpenStudy (kelhina):

can you help me with one more?

OpenStudy (jackthegreatest):

i have 5 min haha sure

OpenStudy (jackthegreatest):

repost...

OpenStudy (kelhina):

@StudyGurl14

OpenStudy (kelhina):

What is the simplified form of this expression, written in scientific notation? (3.276×10^5)×400,000÷(3.9×10^6)

OpenStudy (kelhina):

can you help me @StudyGurl14

OpenStudy (studygurl14):

can you post this in a new thread? this thread is long and my computer is slow :/